OpenStudy (anonymous):

A swimming pool charges $6 for adults and $4 for children. On a particular day, they made $1326 in admissions. If a total of 289 people paid to swim, there must have been ____ adults and _____ children.

OpenStudy (compassionate):

Adults + Children = 1.326 6a + 4c = 1.326 This equation will give you the adults and children revenue The total number of people were 289 Adults + Children = 289 So, the system becomes 6a + 4c = 1,326 a + c = 289 now just solve the system.

OpenStudy (anonymous):

How do you solve the system?

OpenStudy (jdoe0001):

hmm if you solve say.... a + c = 289 for say "c" what do you get?

OpenStudy (anonymous):

I have no idea. I don't know how to figure it out.

OpenStudy (jdoe0001):

hmmm you may want to brush up your linear simplication material then firstly

OpenStudy (compassionate):

a + c = 289 i got bad news, Lauren, if you're having trouble solving for a, then you're really going to hate the process of elimination and substitution. a + c = 289. So solve, do the opposite. So a - c + c = 289 - c a = (289 - c) Now plug it into either equation and solve for a, then plug that back into it and completely solve c.